# EmoLoom-2B: Fast Base-Model Screening for Emotion Classification and VAD with Lexicon-Weak Supervision and KV-Off Evaluation

### v1

@@ Introduction
Early neural approaches treated emotion as single- or multi-label classification, later extending to joint prediction of discrete categories and continuous Valence–Arousal–Dominance (VAD). Recent small language models (SLMs, $\sim$2B) inherit strong text priors and can be instruction-tuned for the joint task, but evaluation often confounds modeling quality with formatting or decoding choices. Common resources include crowd-labeled dialogue and sentence datasets (e.g., GoEmotions, EmpatheticDialogues, DailyDialog) and have encouraged multi-label metrics (Macro-F1/Precision/Recall) alongside continuous VAD error. Our work follows this joint formulation but emphasizes answer validity via a one-line JSON contract and a ParseOK rate so that improvements are not artifacts of prompt or parser idiosyncrasies (dg:20; rs:19; ls:17; r:80; m:18; wkb:13).
@@
Psychological theories describe emotions through cognitive appraisals such as goal attainment, controllability, certainty, and fairness. Computational uses range from feature engineering and rule systems to text classifiers trained on appraisal cues. Many LLM pipelines now generate lengthy rationales to imitate human explanations, yet rationale length becomes a tuning knob and may not translate to better VAD fidelity. We instead adopt appraisal theory as a training-time constraint: a lightweight external “appraisal-atom” verifier provides semantic guidance without fusing explanations into the generator, preserving simplicity, speed, and controllable failure modes (occ:88; se:85; scherer:01; e:92).
@@
Emotion and affective lexica (e.g., NRC Emotion Lexicon) and affective norm lists (e.g., Warriner’s VAD norms) have long enabled low-cost labels or priors. They are effective for coverage and quick diagnostics but are sensitive to domain shift, polysemy, and context (negation, sarcasm). In EmoLoom-2B, lexica are not treated as fixed labels; they supply weak VAD signals to (i) regularize the model through a VAD-preserving consistency loss that aligns generated text with target VAD and (ii) support polarity stress tests via Valence Flip augmentation. This keeps lexical priors useful while avoiding overcommitment to word-level heuristics (mt:13; wkb:13; m:18; pz:06; wb:10; pl:08; rb:20).
@@
Instruction/SFT pipelines for small models typically vary decoding hyperparameters (temperature, top-$p$, max length) and cache usage across training and evaluation, yielding non-trivial score variance and limited replicability. KV-cache in particular changes compute and sometimes output trajectories across hardware and seeds, complicating fair comparison. We advocate a protocol-true setup: identical JSON I/O and KV-off decoding for both training and evaluation, making results comparable across machines and allowing ablations to measure modeling changes rather than decoding drift (bm:20; dc:19; dy:19; hbd:20; liang:22; ba:21; p:21; bgm:21).
@@
Many deployments need a rapid “screening pass” to choose a base model before heavier training or multimodal fusion. Prior work on efficient benchmarking suggests time- or compute-bounded evaluation to trade breadth for speed, but such protocols are rarely standardized for emotion + VAD. EmoLoom-2B provides a simple, auditable quick-eval path—shared metrics, unified logs/figures, and ETA reporting—so candidate backbones can be compared under the same budget (liang:22; rwg:20).
@@
We report Macro-F1/Precision/Recall, VAD $(1-\mathrm{RMSE})$, and ParseOK. The best overall configuration is 20:80 (dg:20; rs:19).
@@
Utterance-level weak VAD is aggregated by weighted mean over covered tokens (weights $\omega_{j}{=}1$); samples with zero coverage are excluded by the $\mathrm{vad\_conf}$ filter (m:18; wkb:13).
@@
For DailyDialog, we split dialogues into utterances, compute weak VAD via the same aggregation, and map its emotion tags into our label space (unseen tags $\rightarrow$ other). The converted file is exported under a distinct name (e.g., dev_dailydialog_weak.jsonl) to avoid overwriting the main dev set. Under a 1-hour quick-eval regime, the 20:80 model reaches Macro-F1 $\approx 0.307$ and ParseOK $\approx 0.976$ (ls:17).
@@
Training uses bf16 forward with TF32 matmul; AdamW ($\beta_{1}{=}0.9,\ \beta_{2}{=}0.95$, weight decay 0.1), cosine LR with $3\%$ warmup, gradient clipping at 1.0, gradient checkpointing, and the OOM self-healing routine (Sec. Fair KV-off Protocol and Reproducibility). Figure 3 shows smoothed loss; the 20:80 mix converges quickest and to the lowest loss (min $\approx 0.160$), matching its best dev trade-off. Figure 4 visualizes LR schedule and gradient norms, indicating stable updates after the initial warmup (akin to curriculum-style scheduling (bengio:09)).

### v2

@@ Introduction
Early neural approaches treated emotion as single- or multi-label classification, later extending to joint prediction of discrete categories and continuous Valence–Arousal–Dominance (VAD). Recent small language models (SLMs, $\sim$2B) inherit strong text priors and can be instruction-tuned for the joint task, but evaluation often confounds modeling quality with formatting or decoding choices. Common resources include crowd-labeled dialogue and sentence datasets (e.g., GoEmotions, EmpatheticDialogues, DailyDialog) and have encouraged multi-label metrics (Macro-F1/Precision/Recall) alongside continuous VAD error. Our work follows this joint formulation but emphasizes answer validity via a one-line JSON contract and a ParseOK rate so that improvements are not artifacts of prompt or parser idiosyncrasies (Demszky et al. 2020; Rashkin et al. 2019; Li et al. 2017; Russell 1980; Mohammad 2018; Warriner et al. 2013).
@@
Psychological theories describe emotions through cognitive appraisals such as goal attainment, controllability, certainty, and fairness. Computational uses range from feature engineering and rule systems to text classifiers trained on appraisal cues. Many LLM pipelines now generate lengthy rationales to imitate human explanations, yet rationale length becomes a tuning knob and may not translate to better VAD fidelity. We instead adopt appraisal theory as a training-time constraint: a lightweight external “appraisal-atom” verifier provides semantic guidance without fusing explanations into the generator, preserving simplicity, speed, and controllable failure modes (Ortony et al. 1988; Smith and Ellsworth 1985; Scherer 2001; Ekman 1992).
@@
Emotion and affective lexica (e.g., NRC Emotion Lexicon) and affective norm lists (e.g., Warriner’s VAD norms) have long enabled low-cost labels or priors. They are effective for coverage and quick diagnostics but are sensitive to domain shift, polysemy, and context (negation, sarcasm). In EmoLoom-2B, lexica are not treated as fixed labels; they supply weak VAD signals to (i) regularize the model through a VAD-preserving consistency loss that aligns generated text with target VAD and (ii) support polarity stress tests via Valence Flip augmentation. This keeps lexical priors useful while avoiding overcommitment to word-level heuristics (Mohammad and Turney 2013; Warriner et al. 2013; Mohammad 2018; Polanyi and Zaenen 2006; Wiegand et al. 2010; Pang and Lee 2008; Ratner et al. 2020).
@@
Instruction/SFT pipelines for small models typically vary decoding hyperparameters (temperature, top-$p$, max length) and cache usage across training and evaluation, yielding non-trivial score variance and limited replicability. KV-cache in particular changes compute and sometimes output trajectories across hardware and seeds, complicating fair comparison. We advocate a protocol-true setup: identical JSON I/O and KV-off decoding for both training and evaluation, making results comparable across machines and allowing ablations to measure modeling changes rather than decoding drift (Brown et al. 2020; Devlin et al. 2019; Dai et al. 2019; Holtzman et al. 2020; Liang et al. 2022; Belz et al. 2021; Pineau et al. 2021; Bender et al. 2021).
@@
Many deployments need a rapid “screening pass” to choose a base model before heavier training or multimodal fusion. Prior work on efficient benchmarking suggests time- or compute-bounded evaluation to trade breadth for speed, but such protocols are rarely standardized for emotion + VAD. EmoLoom-2B provides a simple, auditable quick-eval path—shared metrics, unified logs/figures, and ETA reporting—so candidate backbones can be compared under the same budget (Liang et al. 2022; Ribeiro et al. 2020).
@@
We report Macro-F1/Precision/Recall, VAD $(1-\mathrm{RMSE})$, and ParseOK. The best overall configuration is 20:80 (Demszky et al. 2020; Rashkin et al. 2019).
@@
Utterance-level weak VAD is aggregated by weighted mean over covered tokens (weights $\omega_{j}{=}1$); samples with zero coverage are excluded by the $\mathrm{vad\_conf}$ filter (Mohammad 2018; Warriner et al. 2013).
@@
For DailyDialog, we split dialogues into utterances, compute weak VAD via the same aggregation, and map its emotion tags into our label space (unseen tags $\rightarrow$ other). The converted file is exported under a distinct name (e.g., dev_dailydialog_weak.jsonl) to avoid overwriting the main dev set. Under a 1-hour quick-eval regime, the 20:80 model reaches Macro-F1 $\approx 0.307$ and ParseOK $\approx 0.976$ (Li et al. 2017).
@@
Training uses bf16 forward with TF32 matmul; AdamW ($\beta_{1}{=}0.9,\ \beta_{2}{=}0.95$, weight decay 0.1), cosine LR with $3\%$ warmup, gradient clipping at 1.0, gradient checkpointing, and the OOM self-healing routine (Sec. Fair KV-off Protocol and Reproducibility). Figure 3 shows smoothed loss; the 20:80 mix converges quickest and to the lowest loss (min $\approx 0.160$), matching its best dev trade-off. Figure 4 visualizes LR schedule and gradient norms, indicating stable updates after the initial warmup (akin to curriculum-style scheduling (Bengio et al. 2009)).